Find the Center of mass of a thin plate of constant density 𝛿 covering the region bounded by the parabola 𝑦 = π‘₯2 and the line 𝑦 = 4.


1
Expert's answer
2022-05-12T18:31:22-0400

Solution

Points of intersection of given curves are solution of equation π‘₯2 = 4 => x1 = a = -2, x2 = b = 2.

For density 𝛿 mass of the plate is

"M=\delta\int_{a}^{b}{\left(4-x^2\right)dx=}\delta\left(4x-\frac{1}{3}x^3\right)\left|\begin{matrix}2\\-2\\\end{matrix}\right.=\delta\left(16-\frac{16}{3}\right)=\delta\frac{32}{3}"

Equations of Moments

"M_x=\delta\int_{-2}^{2}{\frac{1}{2}\left[4^2-\left(x^2\right)^2\right]dx}=\delta\left[8x-\frac{1}{10}x^5\right]\left|\begin{matrix}2\\-2\\\end{matrix}\right.=\delta\left[32-\frac{64}{10}\right]=25.6\delta"

"M_y=\delta\int_{-2}^{2}x\left(4-x^2\right)dx=\delta\left[2x^2-\frac{1}{4}x^4\right]\left|\begin{matrix}2\\-2\\\end{matrix}\right.=0"

Center of Mass Coordinates

"x_C=\frac{M_y}{M}=0"

"y_C=\frac{M_x}{M}=\frac{256\bullet3}{320}=2.4"